What is the bond order of Ne2^+, please show work. Is it stable.
Answer #1

ANSWER:

Ne2+ molecular ion having total 19 electron ,

We know electron configuration of Ne2+ is KK (1σ)2(1σ*)2(2π)4(2σ)2(2π*)4(2σ*)1.

Bond order = No.of bonding electron - No .of anti bonding electron /2

Apply above formulea, we get answer easily,

B.O = 8-7/2

B.O = 1/2 =0.5.

Bond order is 0.5. so compound doesnot exits,unstable one.